A chef is using a mixture of two brands of Italian dressing. The first brand contains 10% vinegar, and the second brand contains 15% vinegar. The chef wants to make 230 milliliters of a dressing that is 14% vinegar. How much of each brand should she use?
Answer by algebrahouse.com(1659)

x = number of milliliters of first brand
230 - x = number of milliliters of second brand

0.1x + 0.15(230 - x) = 0.14(230) {first is 10%, second is 15% and total is 14%}
0.1x + 34.5 - 0.15x = 32.2 {used distributive property and multiplied}
-0.05x + 34.5 = 32.2 {combined like terms}
-0.05x = -2.3 {subtracted 34.5 from each side}
x = 46 {divided each side by -0.05}
230 - x = 184 {substituted 46, in for x, into 230 - x}

46 mL of first brand
184 mL of second brand

For more help from me, visit: www.algebrahouse.com